### v3

# Compact Kähler manifolds with quasi-positive holomorphic sectional curvature

Abstract. In this paper, we establish a “pseudo-effective” version of the holonomy principle for compact Kähler manifolds with nonnegative holomorphic sectional curvature. As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), $M$ is projective and rationally connected. This answers a problem raised by Yang and Matsumura and extends Yau’s conjecture. We also get the structure theorem of a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Furthermore, we show that a compact Kähler manifold admits a Hermitian metric with positive real bisectional curvature is a projective and rationally connected manifold.
AMS Mathematics Subject Classification. 53C07, 58E15
Keywords and phrases. Kähler metric, holomorphic sectional curvature, projective, rational connectedness
1. Introduction
A projective manifold $M$ is called rationally connected if any two points on $M$ can be connected by some rational curve. The rational connectedness is an important and fundamental concept in algebraic geometry, and many people have given the criteria for it ([8, 17, 31, 23, 7, 22]). It is also interesting to give geometric interpretations for rational connectedness. Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along the nonzero $V\in T^{1,0}M$ is defined by
$H(V)=\frac{R(V,\overline{V},V,\overline{V})}{|V|_{\omega}^{4}},$
where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section.
Conjecture 1.1 ([39, Problem 47]).
If a compact complex manifold $M$ admits a Kähler metric $\omega$ with positive holomorphic sectional curvature, then $M$ is projective and rationally connected.
By studying MRC fibration, Heier and Wong ([18]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature. Then Yang ([36, 37]) proved that a Kähler manifold $M$ with positive holomorphic sectional curvature satisfies $H^{p,0}_{\bar{\partial}}(M)=H^{0,2}_{\bar{\partial}}(M)=0$ by introducing the notion of RC-positivity and affirmly confirmed Yau’s conjecture, but it’s seems difficult to apply his method to the case of quasi-positive holomorphic sectional curvature. There are some other interesting curvature conditions toward rational connectedness. L.Ni ([28]) introduced the concept of $k$-Ricci curvature for compact Kähler manifold and generalized Yang’s method to show that a compact Kähler manifold with positive $k$-Ricci curvature is projective and rationally connected. Recently, Li, Zhang and the second author ([24, Theorem 1.14]) established the equivalence between mean curvature (quasi-)positivity and rational connectedness for compact Kähler manifolds.
In [25, 26], Matsumura established the structure theorem for a projective manifold $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration: the universal cover of $X$ splits isometrically as $F\times\mathbb{C}^{n-k}$ for some rationally connected manifold $F$. As a corollary, he generalized Yau’s conjecture for projective manifolds ([26, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected. Following the definition in [20], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
$R(V,\overline{X},Y,\overline{Z})=0$
for all $X,Y,Z\in T_{x_{0}}^{1,0}M$. Motivated by these results, Matsumura and Yang proposed the following natural generalization.
Conjecture 1.2 ([37, Conjecture 1.9], [27, Problem 3.3]).
Let $M$ be a compact Kähler manifold. If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point (or the holomorphic sectional curvature is quasi-positive), then $M$ is projective and rationally connected.
More generally, Matsumura proposed the following extension and suggested to investigate Albanese maps instead of MRC fibrations for compact Kähler manifolds. We refer the readers to [27, Section 3], [26, Section 5] for more detailed discussion.
Problem 1.3 ([27, Problem 3.5]).
Can the structure theorem for projective manifolds with nonnegative holomorphic sectional curvature be generalized to compact Kähler manifolds?
In this paper, we establish the following “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (cf. [8, Theorem 1.5]), which is the first main result of our paper.
Theorem 1.4.
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Assume that there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle, then the following statements holds:
  • (1)
    There exists some nonzero truly flat tangent vectors at any point of $M$ determined by $\mathcal{F}$.
  • (2)
    $\mathcal{F}$ is flat as a line bundle and is invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$; in fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
The proof of Theorem 1.4 is based on an integral inequality associated with $\mathcal{F}$ (Proposition 3.1). As a direct application of Theorem 1.4 (1) and Campana-Demailly-Peternell’s criterion for rational connectedness, we get the second main result of our paper, which confirms Conjecture 1.2 for the Kähler case and extends Yau’s conjecture to the quasi-positive case. Our approach is independent of those of Heier-Wong ([18]) and Matsumura ([26]) for projective manifolds.
Theorem 1.5.
Let $(M,\omega)$ be a compact Kähler manifolds. If the holomorphic sectional curvature holomorphic sectional curvature is nonnegative and $(M,\omega)$ has no nonzero truly flat tangent vector at some point, then $M$ is projective and rationally connected.
Remark 1.6.
Statement (2) of Theorem 1.4 is same as the statement of [8, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved. Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([9]), De Rham’s splitting theorem ([15]) and Berger’s classification ([2]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([8, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy, $Y_{j}$ being Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
Statement (1) of Theorem 1.4 ensures that factors of irreducible holonomy $H=SU(n_{j})$ or $Sp(\frac{n_{k}^{\prime}}{2})$ can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature.
Theorem 1.7.
Let $(M,\omega)$ be a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Then $M$ admits a holomorphic and isometric splitting $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ where $M_{l}$ are rationally connected manifolds with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is Hermitian symmetric of compact type).
If the following splitting theorem for nonnegative holomorphic sectional curvature established, Problem 1.3 will be resolved.
Question 1.8.
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Can the universal cover $\widetilde{M}$ of $M$ splits isometrically as $N\times\mathbb{C}^{k}$ for some compact Kähler manifold $N$?
For the Hermitian category, Yang ([36, Corollary 1.10]) proved that a compact Kähler surface admits a Hermitian metric with positive holomorphic sectional curvature is projective and rationally connected. In [38], Yang-Zheng introduced the concept of real bisectional curvature as in the following way and generalized Wu-Yau’s Theorem ([19, 34, 35, 16]) to the Hermitian case. The underlying reason for the extension is that the Schwarz lemma works the same when the target metric is only Hermitian but has nonpositive real bisectional curvature on a Kähler background. Also, they proposed the following generalization of Yau’s Conjecture 1.1 for the Hermitian case ([38, Conjecture 1.6 (b)]).
Conjecture 1.9.
Let $(M,h)$ be a Hermitian manifold with positive real bisectional curvature. Then M is a projective and rationally connected manifold.
We say that a Hermitian manifold $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
$B_{h}(e,a)=\frac{1}{|a|^{2}}\sum\limits_{i,j=1}^{n}R_{i\bar{i}j\bar{j}}a_{i}a_{j}>0,$
where $R$ is the chern curvature tensor of $h$. The term $B_{h}\geq 0,\ B_{h}<0$ or $B_{h}\leq 0$ are defined similarly. This curvature condition is the same as positive holomorphic sectional curvature when the metric $h$ is Kähler and slightly stronger in the non-Kähler case (cf. [38]). Under the entra assumption that $M$ is Kählerian, Conjcture 1.9 is just a weaker version of Conjecture 1.2. In this special case, Tang ([33, Theorem 1.4]) confirmed the Conjecture 1.9 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$. Since the curvature term of Bochner-type formula in the proof of Proposition 3.1 can be controlled by real bisectional curvature, we confirm Conjecture 1.9 under the additional assumption that $M$ is Kählerian.
Theorem 1.10.
Let $(M,h)$ be a compact Hermitian manifold with positive real bisectional curvature. Then for any holomorphic vector bundle $E$ on $M$, there exists a constant $C_{E}>0$ such that
$H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})=0\ \ \text{for all $p,l\in\mathbb{N}_{+}$ with $p\geq C_{E}l$.}$
In particular, if $M$ is Kählerian, $M$ is a projective and rationally connected manifold.
This paper is organized as follows. In section 2, we recall some basic definitions and facts. In section 3, we give a proof of Theorem 1.4 by establishing an integral inequality. In section 4 , we give a proof of Theorem 1.5 and Theorem 1.7. In section 5, we give a proof of Theorem 1.10.
2. Preliminaries
For reader’s convenience, we gather some basic definitions and useful facts known to experts, which will be referenced later in the text.
2.1. Pseudo-effective line bundle and plurisubharmonic function
We first introduce the concept of pseudo-effective line bundle (see e.g. [11, 14]).
Definition 2.1.
A singular Hermitian metric $h$ on a line bundle $L$ is a metric which is given in any trivialization $\theta:F|_{\Omega}\rightarrow\Omega\times\mathbb{C}$ by
$\|\xi\|_{h}=|\theta(\xi)|e^{-\varphi(x)},\ x\in\Omega,\xi\in L_{x}$
where $\varphi\in L^{1}_{loc}(\Omega)$, called the weight of the metric with respect to the trivialization $\theta$.
Definition 2.2.
Let $L$ be a line bundle on a compact complex manifold $X$. We say that $L$ is pseudo-effective if $L$ can be equipped with a singular Hermitian metric $h$ with $T=\frac{\sqrt{-1}}{2\pi}\Theta_{L,h}=\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ as a current, equivalently, the weight $\varphi\in Psh(\Omega)\cap L_{loc}^{1}(\Omega)$ in any trivalization.
Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [13, Chapter 1 and Chapter 3]).
Lemma 2.3.
Let $u\in Psh(\Omega)\cap L^{1}_{loc}(\Omega)$. If $(\rho_{\epsilon})$ is a family of smoothing kernels, then $u*\rho_{\epsilon}\in Psh(\Omega_{\epsilon})\cap C^{\infty}(\Omega_{\epsilon})$, the family $(u*\rho_{\epsilon})$ is decreasing as $\epsilon\rightarrow 0$ and $\lim_{\epsilon\rightarrow 0}u*\rho_{\epsilon}=u,$ where $\Omega_{\epsilon}=\{x\in\Omega:d(x,\partial\Omega)>\epsilon\}$.
Lemma 2.4.
([3, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
$\lim\limits_{j\rightarrow\infty}v_{j}^{i}=v^{i}\in Psh(\Omega)\cap L^{\infty}_{loc}(\Omega),\ 1\leq i\leq k.$
Then
$\lim\limits_{j\rightarrow\infty}v_{j}^{1}dd^{c}v_{j}^{2}\wedge\cdots\wedge dd^{c}v_{j}^{k}=v^{1}dd^{c}v^{2}\wedge\cdots dd^{c}v^{k},$
where the limits is in the weak topology on the currents of bigdegree $(k,k)$ on of order zero.
2.2. Some criterions for rational connectedness
Based on the uniruledness criterion estabished in ([6]), Campana-Demailly-Peternell establish the following criterions for rational connectedness.
Lemma 2.5 ([8, Theorem 1.1]).
Let $M$ be a projective manifold. The following properties are equivalent.
  1. (a)
    $M$ is rationally connected.
  2. (b)
    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$, $1\leq p\leq n$, $\mathcal{F}$ is not pseudo-effective.
  3. (c)
    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(T^{1,0}M^{*})^{\otimes p})$, $p\geq 1$, $\mathcal{F}$ is not pseudo-effective.
  4. (d)
    For any ample line bundle $A$ on $X$, there exists a constant $C_{A}>0$ such that
    $H^{0}(M,(T^{1,0}M^{*})^{\otimes m}\otimes A^{\otimes k})=0$
    for all $m,k\in\mathbb{N}^{*}$ with $m\geq C_{A}k$.
2.3. Some properties of curvature tensor
Let us recall some algebraic properties about curvature tensor.
Lemma 2.6.
(Royden’s observation, [32, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors. If $S(\xi,\bar{\eta},\zeta,\bar{\omega})$ is a symmetric bihermitian form on some complex vector space $V$ (i.e., $S(\xi,\bar{\eta},\zeta,\omega)=S(\zeta,\bar{\eta},\xi,\bar{\omega})$ and $S(\xi,\bar{\omega},\zeta,\bar{\eta})$), such that for all $\xi\in V$, we have
$S(\xi,\bar{\xi},\xi,\bar{\xi})\leq K|\xi|^{4},$
for some constant $K$, then
$\sum\limits_{\alpha,\beta}S(\xi_{\alpha},\bar{\xi}_{\alpha},\xi_{\beta},\bar{\xi}_{\beta})\leq\frac{1}{2}K[\sum\limits|\xi_{\alpha}|^{4}+(\sum\limits|\xi_{\alpha}|^{2})^{2}].$
Lemma 2.7.
([37, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Let $e_{1}\in T_{x_{0}}^{1,0}M$ be a unit vector which minimizes the holomorphic sectional curvature $H$ of $\omega$ at $x_{0}$, then
$2R(e_{1},\overline{e_{1}},W,\overline{W})\geq(1+|\langle W,\overline{e_{1}}\rangle_{g}|^{2})R(e_{1},\overline{e_{1}},e_{1},\overline{e_{1}})$ (2.1)
for any unit vector $W\in T_{x_{0}}^{1,0}M$.
Lemma 2.8.
([20, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Assume that the holomorphic sectional curvature $H$ of $\omega$ is nonnegative at $x_{0}$. If a nonzero tangent vector $V\in T^{1,0}_{x_{0}}M$ such that $R(V,\overline{V},W,\overline{W})=0$ for any $W\in T^{1,0}_{x_{0}}(M)$, then
$R(V,\overline{X},Y,\overline{Z})=0$ (2.2)
for any $X,Y,Z\in T_{x_{0}}^{1,0}M$, i.e. $V$ is a nonzero truly flat tangent vector at $x_{0}$.
3. Proof of Theorem 1.4
In this subsection, we prove Theorem 1.4. The proof is divided into two parts: establishing an integral inequality and performing a pointwise computation. The method is based on integration by parts, Bochner-type formula and algebraic properties of holomorphic sectional curvature.
Setting.
Let $(M,\omega)$ be a compact Kähler manifold. Assume that $\mathcal{F}$ is an invertible sheaf of $\mathcal{O}(\Lambda^{p,0}M)$ which is pseudo-effective as a line bundle. Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the Čech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$. We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([8]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|_{g}}{|\eta_{\alpha}|_{h}}$. Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current,
$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha})$
is also a closed positive current on $U_{\alpha}$. This shows that $\partial\varphi_{\alpha}$ is $L^{2}$ with respect to the wight $e^{\varphi_{\alpha}}$ and $e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}$ has locally finite measure corfficients on $U_{\alpha}$. Hence
$\partial\psi=e^{\varphi_{\alpha}}(\partial\varphi_{\alpha}|\eta_{\alpha}|^{2}_{g}+\partial|\eta_{\alpha}|^{2}_{g})$
is $L^{2}$-integrable on $M$. We will derive the following integral inequality for $\psi$, which is the critical point of the proof of Theorem 1.4.
Proposition 3.1.
The setting is as above, then
$-\int_{M}|\partial\psi|^{2}_{g}\cdot\omega^{n}\geq\frac{p^{2}}{2}\int_{M}\kappa\psi^{2}\cdot\omega^{n}$ (3.1)
where $\kappa$ is a continuous function such that $\kappa(x)$ which is a lower bound of the holomorphic sectional curvature of $\omega$ at $x$.
3.1. Proof of the integral inequality (3.1)
This subsection is devoted to the proof of Proposition 3.1. Firstly, we will proof the following key identity for forms.
Lemma 3.2.
Let $(M,g)$ be a Hermitian manifold of complex dimension $n$, $\alpha$ be a real (1,1)-form and $\eta$ be a (p,0)-form on $M$. We can define a real semi-positive $(1,1)$-form $\beta$ associated to $\eta$ by
$\beta=\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!}\right),$ (3.2)
where $\operatorname{\Lambda}$ is the dual Lefschetz operator with respect to $\omega$. Then
$\operatorname{tr}_{\omega}\beta=|\eta|^{2}_{g}$ (3.3)
and
$\begin{split}(\sqrt{-1})^{p^{2}}\alpha\wedge\eta\wedge\bar{\eta}\wedge\frac{\omega^{n-p-1}}{(n-p-1)!}=\left[\operatorname{tr}_{\omega}\alpha\cdot|\eta|_{g}^{2}-p\langle\alpha,\beta\rangle_{g}\right]\frac{\omega^{n}}{n!},\end{split}$ (3.4)
where $\omega$ is the fundamental (1,1)-form associated to the Hermitian metric $g$.
Proof.
In local holomorphic coordinates $\{z^{1},\cdots,z^{n}\}$, we have $\alpha=\sqrt{-1}\alpha_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$, $\omega=\sqrt{-1}g_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$ and $\eta=\eta_{I_{p}}dz^{I_{p}}$, where $I_{p}=(i_{1},\cdots,i_{p})\in\mathbb{N}_{+}^{p}$ and $dz^{I_{p}}=dz^{i_{1}}\wedge\cdots\wedge dz^{i_{p}}$. For any $x_{0}\in M$, we can choose local holomorphic coordinates centered at $x_{0}$ such that
$\omega=\sqrt{-1}\sum\limits_{i}dz^{i}\wedge d\bar{z}^{i}$
and
$\alpha=\sqrt{-1}\sum\limits_{i}\alpha_{i\bar{i}}dz^{i}\wedge d\bar{z}^{i}$
at $x_{0}$. So, we have
$\omega^{n-p-1}$
$=$ $(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}},$
and
$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}$
$=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
+(terms involving ηJpηKp¯ with Jpπ(Kp) for some permutation $+(\text{terms involving $\eta_{J_{p}}\overline{\eta_{K_{p}}}$ with $J_{p}\neq\pi(K_{p})$ for some permutation $\pi$}).$
It follows that
$\begin{split}&(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}\wedge\omega^{n-p-1}\\ =&(\sqrt{-1})^{n}n!\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{1}\wedge d\bar{z}^{1}\wedge\cdots\wedge dz^{n}\wedge d\bar{z}^{n}\\ =&\left(\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}\right)\omega^{n}\\ =&\left(\sum\limits_{i,J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}-\sum\limits_{i\in J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}\right)\omega^{n}\\ =&\left(\sum\limits_{i,J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}-p\sum\limits_{i,I_{p-1}}\alpha_{i\bar{i}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\right)\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|_{g}^{2}}{p!}-\langle\alpha\wedge\omega^{p-1},(\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!(p-1)!}\rangle_{g}\right]\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|^{2}_{g}}{p!}-\frac{1}{(p-1)!}\langle\alpha,\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!}\right)\rangle_{g}\right]\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|^{2}_{g}}{p!}-\frac{1}{(p-1)!}\langle\alpha,\beta\rangle_{g}\right]\omega^{n}.\end{split}$ (3.5)
This verifies (3.4). Locally, $\beta=\sqrt{-1}\beta_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ and
$\beta_{i\bar{j}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{jI_{p-1}}}.$ (3.6)
We can reselect the local holomorphic coordiates such that $(g_{i\bar{j}})$ is still the identity and $(\beta_{i\bar{j}})$ is diagonal at $x_{0}$. Then
$\beta_{i\bar{i}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\geq 0.$ (3.7)
Hence $\beta$ is semi-positive and
$\operatorname{tr}_{\omega}\beta=p!\sum\limits_{i,I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}=|\eta|^{2}_{g}.$ (3.8)
The proof is completed. ∎
Proof of Proposition 3.1.
Now let us proceed to show Proposition 3.1. It follows from Lemma 3.2 that
$\begin{split}n&\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}=e^{\varphi_{\alpha}}\operatorname{tr}_{\omega}(\sqrt{-1}\partial\bar{\partial}\psi)|\eta_{\alpha}|^{2}_{g}\cdot\omega^{n}\\ =&p\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}\end{split}$ (3.9)
where $\beta_{\alpha}=\Lambda^{p-1}(\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{(p-1)!})$. $e^{\varphi_{\alpha}}\beta_{\alpha}$ and $e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}$ are globally defined differential forms with bounded measurable cofficients. For clarity, the arguement will be divided into three steps.
Step 1. The integration by parts
Although $\varphi_{\alpha}$ is in general singular, we can still do the integration by parts as in [12]. For reader’s convenience, we give a detailed explaination. Take a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$. Since $\mathrm{supp}\rho_{\alpha}\subset\subset U_{\alpha}$, for any $\epsilon>0$ sufficiently small, there exists a sequence of smooth plurisubharmonic functions $\varphi_{\alpha,\epsilon}$ decreasing to $\varphi_{\alpha}$ on some open subset $V_{\alpha}$ such that $\mathrm{supp}\rho_{\alpha}\subset\subset V_{\alpha}\subset\subset U_{\alpha}$. Now the results of Bedford-Taylor (Lemma 2.4) applied to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, imply that we have local weak covergence
$e^{\varphi_{\alpha,\epsilon}}\partial\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha},\ e^{\varphi_{\alpha,\epsilon}}\rightarrow e^{\varphi_{\alpha}}\ \text{and}\ e^{\varphi_{\alpha,\epsilon}}\partial\varphi_{\alpha,\epsilon}\wedge\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}$
on $V_{\alpha}$, then
$\begin{split}&-\int_{M}\left|\partial\psi\right|^{2}_{g}\cdot\omega^{n}=-n\int_{M}\partial\psi\wedge\bar{\partial}\psi\wedge\sum\limits_{\alpha}\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\sum\limits_{\alpha}\int_{V_{\alpha}}e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\int_{M}\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}.\end{split}$ (3.10)
The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combine with (3.9), we get
$\begin{split}0\geq&-\int_{M}|\partial\psi|_{g}^{2}\cdot\omega^{n}\\ =&p\int_{M}\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\int_{M}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}.\end{split}$ (3.11)
Next, let us separately analysis each term on the right-hand side of integral inequality (3.11). Firstly, we consider the sign of first term of (3.11).
Step 2. The Bochner formula.
For any $x_{0}\in U_{\alpha}$, we choose local holomorphic coordinates centered at $x_{0}$ such that
$\omega=\sqrt{-1}\sum\limits_{i}dz^{i}\wedge d\bar{z}^{i}$
and
$\beta_{\alpha}=\sqrt{-1}\sum\limits_{i}\beta_{i\bar{i}}dz^{i}\wedge d\bar{z}^{i}$
at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$, recall (3.3) and (3.7) that we have
$\operatorname{tr}_{\omega}\beta=|\eta|_{g}^{2}$ (3.12)
and
$\beta_{i\bar{i}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\geq 0.$ (3.13)
By a direct calculation, we derive the following Bochner type formula in the sense of distribution,
$\begin{split}\partial_{u}\partial_{\bar{u}}(e^{\varphi}|\eta|^{2}_{g})=&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p!\sum\limits_{I_{p}}\sum\limits_{k=1}^{p}\sum\limits_{l=1}^{n}R_{u\bar{u}i_{k}\bar{l}}\eta_{I_{p}}\overline{\eta_{i_{1}\cdots(l)_{k}\cdots i_{p}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p(p!)\sum\limits_{l,k=1}^{n}R_{u\bar{u}l\bar{k}}\sum\limits_{J_{p-1}}\eta_{lJ_{p-1}}\overline{\eta_{kJ_{p-1}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p\sum\limits_{l,k}R_{u\bar{u}l\bar{k}}\beta_{l\bar{k}})\end{split}$ (3.14)
for all $u\in T^{1,0}_{x_{0}}M$, where $R_{i\bar{j}k\bar{l}}=R(\partial_{i},\overline{\partial_{j}},\partial_{k},\overline{\partial_{l}})$ and $D^{\prime}$ is the $(1,0)$-part of the connection $D$ of $\Lambda^{p,0}M$ inducced by Chern connection of $(T^{1,0}M,\omega)$. The second equality follows from interchanging the order of summation. Hence
$\begin{split}&p\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}=e^{2\varphi}(p|\eta|^{2}\partial_{i}\partial_{\bar{i}}\varphi\beta_{i\bar{i}}+p\sum\limits_{i}|D^{\prime}_{i}\eta+\eta\partial_{i}\varphi|^{2}\beta_{i\bar{i}}+p^{2}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}})\\ =&e^{2\varphi}\left(p\psi\langle\sqrt{-1}\partial\bar{\partial}\varphi,\beta\rangle+p\left<\sqrt{-1}\langle D^{\prime}\eta+\eta\partial\varphi,\overline{D^{\prime}\eta+\eta\partial\varphi}\rangle_{g},\beta\right>_{g}+p^{2}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}\right).\end{split}$ (3.15)
It follows from Royden’s observation(Lemma 2.6) that
$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=&e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\\ \geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
at $x_{0}$. The fact that $\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ and $\beta\geq 0$ implies that
$e^{2\varphi}(p\psi\langle\partial\bar{\partial}\varphi,\beta\rangle+p\left<\sqrt{-1}\langle D^{\prime}\eta+\eta\partial\varphi,\overline{D^{\prime}\eta+\eta\partial\varphi}\rangle_{g},\beta\right>_{g}\geq 0.$ (3.17)
Hence, we have
$p\int_{M}\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}\cdot\omega^{n}=p\int_{M}\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}\cdot\omega^{n}\geq\frac{p^{2}}{2}\int_{M}\kappa\psi^{2}\cdot\omega^{n}.$
Step 3. The sign of the second term of (3.11):
As
¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
$=$ $(\sqrt{-1})^{2p}e^{\varphi_{\alpha}}(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})\wedge\overline{(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})}+e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}},$
by applying the integration by parts which is valid as previously explained, we get
$(\sqrt{-1})^{p^{2}}\int_{M}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\wedge\omega^{n-p-1}$
$=$ $(\sqrt{-1})^{p^{2}}\int_{M}\psi\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta}_{\alpha}\wedge\omega^{n-p-1})$
$=$ $(\sqrt{-1})^{p^{2}}\int_{M}\psi\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta}_{\alpha})\wedge\omega^{n-p-1}$
$=$ $\int_{M}\psi(\sqrt{-1})^{(p+1)^{2}}e^{\varphi_{\alpha}}\gamma_{\alpha}\wedge\overline{\gamma_{\alpha}}\wedge\omega^{n-p-1}+\int_{M}\psi\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}\wedge e^{\varphi_{\alpha}}(\sqrt{-1})^{p^{2}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\wedge\omega^{n-p-1}$
$\geq$ $0$
where $\gamma_{\alpha}=\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha}$. The last inequality follows from the fact
$\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}\geq 0,\ (\sqrt{-1})^{(p+1)^{2}}\gamma_{\alpha}\wedge\overline{\gamma_{\alpha}}\geq 0\text{ and }(\sqrt{-1})^{p^{2}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\geq 0.$
In summary, we get the desired integral inequality. The proof is completed. ∎
3.2. The proof of Theorem 1.4
In this subsection, we proceed to show Theorem 1.4. The method is based on performing a pointwise computation.
Proof of Theorem 1.4.
Let $\mathcal{F}$ be an invertible subsheaf of $\mathcal{O}({\Lambda^{p,0}M})$ that is pseudo-effective. Since the holomorphic sectional curvature of $\omega$ is nonnegative on $M$, we can choose $\kappa\equiv 0$ in the setting 3. By Proposition 3.1, we have
$-\int_{M}|\partial\psi|_{g}^{2}\cdot\omega^{n}\geq 0.$
Hence $d\psi=0$ almost everywhere on $M$, which implies that $\psi=C$ for some constant $C\geq 0$ almost everywhere. We claim that $C>0$. Suppose not, we have $\psi=0$ almost everywhere on $U_{\alpha}$. Since $\varphi_{\alpha}\in Psh(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$, $\{\varphi_{\alpha}=-\infty\}$ has zero Lebesgue measure. This fact implies that $\eta_{\alpha}=0$ almost everywhere on $U_{\alpha}$. Then $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$. Therefore, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$. The following calculation will be done in the smooth sense.
For any $\alpha$, take an arbitrary point $x_{0}$ of $U_{\alpha}$. For simplicity, we omit the index $\alpha$. Recall that we can define a real semi-positive $(1,1)$-form $\beta$ associated to $\eta$ by
$\beta=\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\overline{\eta}}{p!}\right)$ (3.18)
such that
$\operatorname{tr}_{\omega}\beta=|\eta|^{2}_{g}>0.$ (3.19)
We choose local holomorphic coordinates centered at $x_{0}$ such that $(g_{i\bar{j}})$ is the identity and $(\beta_{i\bar{j}})$ is diagonal at $x_{0}$. Since $\beta_{i\bar{i}}$ are nonnegative for all $1\leq i\leq n$ and $\sum\limits_{i}\beta_{i\bar{i}}>0$, without loss of generality, we may assume that
$\beta_{1\bar{1}}\geq\beta_{2\bar{2}}\geq\cdots\geq\beta_{m\bar{m}}>0=\beta_{m+1\overline{m+1}}=\cdots=\beta_{n\bar{n}}$ (3.20)
for some positive integer $m\leq n$. Denote
[…]
From the Bochner type formula (3.14), we have
$0=e^{-\varphi}\partial_{v}\partial_{\bar{v}}(e^{\varphi}|\eta|^{2}_{g})=|\eta|^{2}_{g}\partial_{v}\partial_{\bar{v}}\varphi+|D^{\prime}_{v}\eta+\eta\partial_{v}\varphi|^{2}+p\sum\limits_{j=1}^{m}R_{v\bar{v}j\bar{j}}\beta_{j\bar{j}}\geq p\sum\limits_{j=1}^{m}R_{v\bar{v}j\bar{j}}\beta_{j\bar{j}}\lambda_{j}^{2}$ (3.21)
for all $v\in T_{x_{0}}^{1,0}M$. Let $t=(t_{1},\cdots,t_{m})$ be the complex coordinates of $\mathbb{C}^{m}$ and $d\theta$ be the standard volume measure of the unit sphere $S_{2m-1}:=\{t\in\mathbb{C}^{m}:\sum\limits_{i=1}^{m}|t_{i}|^{2}=1\}$. By the Berger’s averaging trick ([2]), we get
[…]
For any $t\in S_{2m-1}$,
$\sum\limits_{i,j,k,l=1}^{m}R_{i\bar{j}k\bar{l}}\lambda_{i}t_{i}\overline{\lambda_{j}}\overline{t_{j}}\lambda_{k}t_{k}\overline{\lambda_{l}}\overline{t_{l}}=R(V_{t},\overline{V_{t}},V_{t},\overline{V_{t}})\geq 0,$ (3.23)
where $V_{t}:=\sum\limits_{i=1}^{m}\lambda_{i}t_{i}\frac{\partial}{\partial z^{i}}$. Combining (3.22) with (3.23), $R(V_{t},\overline{V_{t}},V_{t},\overline{V_{t}})=0,\ \forall t\in S_{2m-1}$. Taking $t_{0}\in S_{2m-1}$ such that only the $l$-th component is $1$ and all other components are 0, then $V_{t_{0}}=\lambda_{l}\frac{\partial}{\partial z^{l}}$ and
$R_{l\bar{l}l\bar{l}}=\frac{1}{\lambda_{l}^{4}}R(V_{t_{0}},\overline{V_{t_{0}}},V_{t_{0}},\overline{V_{t_{0}}})=0,\ \forall 1\leq l\leq m.$ (3.24)
Since the holomorphic sectional curvature is nonnegative, $\frac{\partial}{\partial z^{l}}$ minimizes the holomorphic sectional curvature of $\omega$ at $x_{0}$. It follows from Lemma 2.7 that
$R_{l\bar{l}v\bar{v}}\geq 0,\ \forall 1\leq l\leq m,v\in T^{1,0}_{x_{0}}M.$
Since $\partial_{v}\partial_{\bar{v}}\varphi\geq 0$, (3.21) implies that $R_{l\bar{l}v\bar{v}}=0,\ \forall 1\leq l\leq m$ and
$\partial_{v}\partial_{\bar{v}}\varphi=0,\ D^{\prime}_{v}\eta+\eta\partial_{v}\varphi=0$
[…]
Since $x_{0}\in M$ is arbitrary, we have ¯$\partial\bar{\partial}\varphi=0$ on $M$ and $D^{\prime}\eta+\eta\partial\varphi=0$ on $M$. ¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle. In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, then $\varphi_{\alpha}$ can be written $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$ and
$D^{\prime}(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$
where $e^{f_{\alpha}}\eta_{\alpha}$ is a local holomorphic section of $\mathcal{F}|_{U_{\alpha}}$. This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$. The proof is completed. ∎
4. Proof of Theorem 1.5 and Theorem 1.7
In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7. The proof of Theorem 1.5 is based on Theorem 1.4 (1) and Campana-Demailly-Peternell’s critetion. The proof of Theorem 1.7 follows the arguement in [7, section 4] for $\operatorname{Ric}\geq 0$, demonstrating that factors of irreducible holonomy $H=U(n)$ are rationally connected manifolds. The difference is that factors of irreducible holonomy $H=SU(n)$ or $Sp(\frac{n}{2})$ can be excluded by statement (1) of Theorem 1.4.
Proof of Theorem 1.5.
Since the holomorphic sectional curvature is nonnegative and there exists no nonzero truly flat tangent vector, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective by the statement (1) of Theorem 1.4. This fact implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$, in particular, M is projective by Kodaira’s theorem ([21]). Combine with Campana-Demailly-Peternell’s criterion (Lemma 2.5), $M$ is rationally connected. ∎
Proof of Theorem 1.7.
Let $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ be the De Rham decomposition of $(M,\omega)$, induced by a decomposition of the holonomy representation in irreducible representations ([4]). Since the holonomy is contained in $U(n)$, all factors $(M_{l},\omega_{l})$ are Kähler manifolds with irreducible holonomy and holonomy group $H_{l}\subset U(n_{l})$, $n_{l}=\dim M_{l}$. The product structure implies that $\omega_{l}$ has nonnegative holomorphic sectional curvature. By Berger’s classification of holonomy groups ([4]), there are only three possibilities: $H_{l}=U(n_{l}),\ SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$, unless $M_{l}$ is a Hermitian symmetric space, then necessarily of compact type (such symmetric space were classified by Élie Cartan, they are rational homogeneous, hence rationally connected, their holonomy groups are also known, see. e.g. [5]).
Case 1: $H_{l}=SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$. Then $M_{l}$ is a Calabi-Yau manifold or a holomorphic symplectic manifold (see e.g. [1]). Recall that a simply connected Calabi-Yau manifold $M_{l}$ is a compact simply connected projective manifold with $K_{M_{l}}\simeq O_{M_{l}}$ and $H^{0}(M_{l},\Omega_{M_{l}}^{p})=0$ for $0<p<\dim M_{l}$. A holomorphic symplectic manifold is a compact Kähler manifold $X$ which admits a holomorphic symplectic $(2,0)$-form (of maximal rank everywhere). In particular, $M_{l}$ admits a holomorphic $(n_{l},0)$-form that vanishes nowhere. Recall that we can define a nonnegative $(1,1)$-form $\beta=\Lambda^{p-1}({\frac{\eta\wedge\overline{\eta}}{n_{l}!}})$. For any point $x_{0}$ of $M_{l}$, we can choose some local holomorphic coordinates such that the coefficients of $w_{l}$ is identity and $\beta$ is diagonal at $x_{0}$. Locally, we can write $\eta=\frac{1}{n_{l}!}\sum\limits fdz^{1}\wedge\cdots dz^{n}$. It follows from the local representation (3.7) of $\beta$ that we have
$\beta=\frac{|\eta|^{2}}{n_{l}}\sum\limits_{i=1}^{n_{l}}dz^{i}\wedge d\overline{z}^{i}$
at $x_{0}$. The proof of Theorem 1.4 tells us that $\{\frac{\partial}{\partial z_{i}}\}_{i=1}^{n_{l}}$ are truly flat tangent vectors at $x_{0}$. Hence the Riemannian curvature tensor $\operatorname{Rm}$ of $\omega_{l}$ vanishes at $x_{0}$. Since $x_{0}$ is arbitrary, $\operatorname{Rm}$ vanishes everywhere, $M_{l}$ is a simply connected complete manifold with zero holomorphic sectional curvature. By classification of complex space forms, $M_{l}$ must be holomorphically isometric to complex Euclidean space $\mathbb{C}^{n_{l}}$, which contradicts the fact that $M_{l}$ is compact.
Case 2: $H_{l}=U(n_{l})$. Since the natural representation of $U(n_{l})$ on $(\mathbb{C}^{n_{l}})^{\otimes p}$ has no invariant line on which $U(n_{l})$ induces a trivial action, there cannot exist any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M_{i})$, $1\leq p\leq n$ such that $\mathcal{F}$ is invariant by parallel transport induced by the connection $D$. Combine with statement (2) of Theorem 1.4, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective. Therefore $M_{l}$ is rationally connected by Campana-Demailly-Peternell’s criterion.
In summary, each $(M_{l},\omega_{l})$ is a rationally connected manifold with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is a hermitian symmetric space of compact type). ∎
5. Proof of Theorem 1.10
In this section, we give a proof of Theorem 1.10.
Proof of Theorem 1.10.
Denote $n=\dim M$ and $r=\operatorname{rank}E$. For any $t\in\mathbb{N}_{+},\ m\in[0,k]\bigcap\mathbb{N}$ and $I_{k}=(i_{1},\cdots,i_{k})\in\mathbb{N}_{+}^{k}$ for some $k\in\mathbb{N}$, we define
$I_{k}(m,t)=(i_{1},\cdots,i_{m-1},t,i_{m},\cdots,i_{k})\in\mathbb{N}_{+}^{k+1}.$
We equip $E$ with a Hermitian metric $H$ and denote its chern curvature tensor by $R^{H}$. Set
$\delta_{2}=\min\limits_{x\in M}\min_{u\in T^{1,0}_{x}M\setminus 0,V\in E_{x}\setminus 0}-\frac{R^{H}(u,\bar{u},V,\bar{V})}{|u|^{2}_{h}|V|^{2}_{H}}.$ (5.1)
Since $(M,h)$ has positive real bisectional curvature, there exists $\delta_{1}>0$ such that for any $x\in M$, any unitary basis $\{\frac{\partial}{\partial z^{1}},\cdots,\frac{\partial}{\partial z^{n}}\}$ of $T_{x}^{1.0}M$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, we have that
$\sum\limits_{i,j=1}^{n}R_{i\bar{i}j\bar{j}}a_{i}a_{j}>\delta_{1}|a|^{2}.$ (5.2)
For any $x\in M$, we choose local holomorphic coordinates $(z^{1},\cdots,z^{n})$ centered at $x$ such that $h_{i\bar{j}}$ is the identity at $x$. Let $\psi$ be any holomorphic section of $H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$. Locally
$\psi=\sum\limits_{I_{p},J_{l}}a_{I_{p},J_{l}}dz^{i_{1}}\otimes\cdots\otimes dz^{i_{p}}\otimes e^{j_{1}}\otimes\cdots\otimes e^{j_{l}}$
with $I_{p}=(i_{1},\cdots,i_{p})\in\mathbb{N}_{+}^{p},J_{l}=(j_{1},\cdots,j_{l})\in\mathbb{N}_{+}^{l}$, where $\{e^{k}\}_{k=1}^{r}$ is a local holomorphic basis of $E$ such that $(H_{t\bar{s}})$ is diagonal at $x$. By a direct computation, we have the following Bochner type formula,
$\partial_{u}\partial_{\bar{u}}|\psi|^{2}$
$=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
$-\sum\limits_{t,s=1}^{r}R^{H}_{u\bar{u}t\bar{s}}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,s)}}$
holds for all $u\in T_{x}^{1,0}M$. Define a real nonnegative $(1,1)$-form $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ by
$\alpha_{i\bar{j}}=\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}=\sum\limits_{m=1}^{p}\langle\iota^{m}_{\frac{\partial}{\partial z^{i}}}(\psi),\overline{\iota^{m}_{\frac{\partial}{\partial z^{j}}}(\psi)}\rangle$ (5.3)
with $\iota^{m}_{\frac{\partial}{\partial z^{i}}}(\psi)=\psi(\cdots,\frac{\partial}{\partial z^{i}},\cdots;\cdots)$ where the contractions are done at the $m$-th slot. In a similar way, define a section $\beta=\beta_{t\bar{s}}e^{t}\otimes\overline{e^{s}}$ of $E\otimes\overline{E}$ such that the components $(\beta_{t\bar{s}})$ is a nonnegative Hermitian matrix by
$\beta_{t\bar{s}}=\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,s)}}=\sum\limits_{k=1}^{l}\langle\iota^{p+k}_{(e^{t})^{*}}(\psi),\overline{\iota^{p+k}_{(e^{s})^{*}}(\psi)}\rangle$ (5.4)
with $\iota^{p+k}_{(e^{t})^{*}}(\psi)=\psi(\cdots;\cdots,(e^{t})^{*},\cdots)$ where the contractions are done at the $(p+k)$-th slot. Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then
$\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=\sum\limits_{i=1}^{n}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,i),J_{l}}}=p|\psi|^{2}$ (5.5)
and
$\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=\sum\limits_{t=1}^{r}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,t)}}=l|\psi|^{2}.$ (5.6)
So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,
$\langle\sqrt{-1}\partial\bar{\partial}|\psi|^{2},\alpha\rangle_{h}=$ $\sum\limits_{i=1}^{n}\partial_{i}\partial_{\bar{i}}|\psi|^{2}\alpha_{i\bar{i}}$
$\geq$ $\sum\limits_{i,j=1}^{n}R^{h}_{i\bar{i}j\bar{j}}\alpha_{i\bar{i}}\alpha_{j\bar{j}}-\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{r}R^{H}_{i\bar{i}t\bar{t}}\alpha_{i\bar{i}}\beta_{t\bar{t}}$
$\geq$ $\delta_{1}\sum\limits_{i=1}^{n}|\alpha_{i\bar{i}}|^{2}+\delta_{2}\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}\sum\limits_{t=1}^{r}\beta_{t\bar{t}}$
$\geq$ $\delta_{1}\frac{(\sum\limits_{i=1}^{n}\alpha_{i\bar{i}})^{2}}{n}+\delta_{2}(\sum\limits_{i=1}^{n}\alpha_{i\bar{i}})(\sum\limits_{t=1}^{r}\beta_{t\bar{t}})$
$=$ $(p+\frac{n\delta_{2}}{\delta_{1}}l)\frac{p\delta_{1}|\psi|^{4}}{n}$
holds for any $x\in M$. Let $C_{E}=\max\{1,-\frac{n\delta_{2}}{\delta_{1}}+2\}$. We assume that $|\psi|^{2}$ achieves its maximum at some point $x_{0}\in M$. If $p,l\in\mathbb{N}_{+}$ with $p\geq C_{E}l$ and $\psi\neq 0$, then we have
$00\geq\langle\sqrt{-1}\partial\bar{\partial}|\psi|^{2},\alpha\rangle_{h}\geq(p+\frac{n\delta_{2}}{\delta_{1}})|\psi|^{4}\geq 2l|\psi|^{4}>$ 0
at $x_{0}$, which is a contradiction. Hence
$H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})=0\ \text{ for any }p\geq C_{E}l.$ (5.7)
The vanishing theorem (5.7) implies that $H^{2,0}_{\bar{\partial}}(M)=0$. In particular, if $M$ is Kählerian, $M$ is projective by Kodaira’s theorem ([21]). Then (5.7) and Campana-Peternell-Demailly’s criterion (Lemma 2.5) imply that $M$ is rationally connected. The proof is completed. ∎

### v4

@@
-Compact Kähler manifolds with quasi-positive holomorphic sectional curvature
+On the structure of compact Kähler manifolds with nonnegative holomorphic sectional curvature
 Abstract.
 In this paper, we establish a “pseudo-effective” version of the holonomy principle for compact Kähler manifolds with nonnegative holomorphic sectional curvature.
-As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), $M$ is projective and rationally connected.
-This answers a problem raised by Yang and Matsumura and extends Yau’s conjecture.
-We also get the structure theorem of a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature.
+As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ must be projective and rationally connected.
+This result answers a problem raised by Yang and Matsumura and extends the Yau’s conjecture.
+We also prove that a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature is projective and rationally connected.
+Additionally, we classify non-projective compact Kähler 3-dimensional manifolds.
 Furthermore, we show that a compact Kähler manifold admits a Hermitian metric with positive real bisectional curvature is a projective and rationally connected manifold.
@@
 The rational connectedness is an important and fundamental concept in algebraic geometry, and many people have given the criteria for it ([9, 19, 37, 27, 8, 26]).
-It is also interesting to give geometric interpretations for rational connectedness.
-Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along the nonzero $V\in T^{1,0}M$ is defined by
+It is also important to provide geometric interpretations for rational connectedness.
+Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along a nonzero $V\in T^{1,0}M$ is defined by
 $H(V)=\frac{R(V,\overline{V},V,\overline{V})}{|V|_{\omega}^{4}},$
-where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section.
-Conjecture 1.1 ([39, Problem 47]).
+where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section:
+Conjecture 1.1 ([45, Problem 47]).
 If a compact complex manifold $M$ admits a Kähler metric $\omega$ with positive holomorphic sectional curvature, then $M$ is projective and rationally connected.
-By studying MRC fibration, Heier and Wong ([18]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature.
-Then Yang ([36, 37]) proved that a Kähler manifold $M$ with positive holomorphic sectional curvature satisfies $H^{p,0}_{\bar{\partial}}(M)=H^{0,2}_{\bar{\partial}}(M)=0$ by introducing the notion of RC-positivity and affirmly confirmed Yau’s conjecture, but it’s seems difficult to apply his method to the case of quasi-positive holomorphic sectional curvature.
-There are some other interesting curvature conditions toward rational connectedness.
-L.Ni ([28]) introduced the concept of $k$-Ricci curvature for compact Kähler manifold and generalized Yang’s method to show that a compact Kähler manifold with positive $k$-Ricci curvature is projective and rationally connected.
-Recently, Li, Zhang and the second author ([24, Theorem 1.14]) established the equivalence between mean curvature (quasi-)positivity and rational connectedness for compact Kähler manifolds.
-In [25, 26], Matsumura established the structure theorem for a projective manifold $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration: the universal cover of $X$ splits isometrically as $F\times\mathbb{C}^{n-k}$ for some rationally connected manifold $F$.
-As a corollary, he generalized Yau’s conjecture for projective manifolds ([26, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected.
-Following the definition in [20], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
-$R(V,\overline{X},Y,\overline{Z})=0$
-for all $X,Y,Z\in T_{x_{0}}^{1,0}M$. Motivated by these results, Matsumura and Yang proposed the following natural generalization.
-Conjecture 1.2 ([37, Conjecture 1.9], [27, Problem 3.3]).
+Heier and Wong ([21]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature by studying MRC fibration.
+By introducing the notion of RC-positivity and studying vanishing theorems, Yang ([42, 43]) showed that a Kähler manifold $M$ with positive holomorphic sectional curvature is projective and rationally connected, then affirmly confirmed Yau’s conjecture.
+There are some other interesting curvature conditions related to rational connectedness (see, for instance, [34, 36, 42, 43]).
+Recently, Li, Zhang and the second author ([29, Theorem 1.14]) established the equivalence between mean curvature quasi-positivity of the tangent bundle and rational connectedness for compact Kähler manifolds.
+In [30, 31], Matsumura established a structure theorem for projective manifolds $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration.
+Both the methods of Heier-Wong and Matsumura rely on the uniruledness criterion for projective manifolds ([7]).
+According to Matsumura’s structure theorem, the universal cover of $X$ splits isometrically and holomorphically as $F\times\mathbb{C}^{n-k}$, where $F$ is a rationally connected manifold.
+This result generalizes the structure theorem established by Howard, Smith and Wu ([24]) and Mok ([33]) for nonnegative holomorphic bisectional curvature.
+As a corollary, Matsumura generalized Yau’s conjecture for projective manifolds ([31, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected.
+According to [23], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
+$R(V,\overline{X},Y,\overline{Z})=0,\text{\ for all $X,Y,Z\in T^{1,0}M$}.$
+Motivated by these results, Matsumura and Yang proposed the following natural generalization.
+Conjecture 1.2 ([43, Conjecture 1.9], [32, Problem 3.3]).
 Let $M$ be a compact Kähler manifold.
-If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point (or the holomorphic sectional curvature is quasi-positive), then $M$ is projective and rationally connected.
-More generally, Matsumura proposed the following extension and suggested to investigate Albanese maps instead of MRC fibrations for compact Kähler manifolds.
-We refer the readers to [27, Section 3], [26, Section 5] for more detailed discussion.
-Problem 1.3 ([27, Problem 3.5]).
+If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point, then $M$ is projective and rationally connected.
+More generally, Matsumura proposed the following extension and suggested investigating Albanese maps instead of MRC fibrations for compact Kähler manifolds, due to the lack of a uniruledness criterion in the Kähler case.
+We refer the reader to [32, Section 3] and [31, Section 5] for a more detailed discussion.
+Problem 1.3 ([32, Problem 3.5]).
 Can the structure theorem for projective manifolds with nonnegative holomorphic sectional curvature be generalized to compact Kähler manifolds?
-In this paper, we establish the following “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (cf.
-[8, Theorem 1.5]), which is the first main result of our paper.
+Towards Conjecture 1.2 and Problem 1.3, we establish a “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (c.f.
+[9, Theorem 1.5] for nonnegative Ricci curvature).
+This is the first main result of our paper.
 Theorem 1.4.
 Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature.
-Assume that there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle, then the following statements holds:
+Assume there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle.
+Then the following statements holds:
   • (1)
-    There exists some nonzero truly flat tangent vectors at any point of $M$ determined by $\mathcal{F}$.
+    $\mathcal{F}$ is flat as a line bundle and invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$.
+In fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
   • (2)
-    $\mathcal{F}$ is flat as a line bundle and is invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$; in fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
-The proof of Theorem 1.4 is based on an integral inequality associated with $\mathcal{F}$ (Proposition 3.1).
-As a direct application of Theorem 1.4 (1) and Campana-Demailly-Peternell’s criterion for rational connectedness, we get the second main result of our paper, which confirms Conjecture 1.2 for the Kähler case and extends Yau’s conjecture to the quasi-positive case.
-Our approach is independent of those of Heier-Wong ([18]) and Matsumura ([26]) for projective manifolds.
+    $\mathcal{F}$ determines a parallel and self-adjoint linear tranformation $S:T^{1,0}M\rightarrow T^{1,0}M$, which induces a decomposition $V+W$ of $T^{1,0}M$ where $V$ and $W$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S$, respectively; in fact, $\forall p\in M$, every $v\in V_{p}$ is truly flat and $\dim V_{p}=\operatorname{rank}(S)>0$.
+The core of the proof of Theorem 1.4 is an integral inequality associated with $\mathcal{F}$ (Proposition 3.1).
+As a direct application of Theorem 1.4 and the Campana-Demailly-Peternell’s criterion for rational connectedness (2.5), we obtain the second main result of our paper, which confirms Conjecture 1.2 for Kähler manifolds and extends Yau’s conjecture to the quasi-positive case.
+Our approach is independent of those by Heier-Wong ([21]) and Matsumura ([31]) for projective manifolds.
 Theorem 1.5.
@@
 Remark 1.6.
-Statement (2) of Theorem 1.4 is same as the statement of [8, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved.
-Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([9]), De Rham’s splitting theorem ([15]) and Berger’s classification ([2]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([8, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy, $Y_{j}$ being Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
-Statement (1) of Theorem 1.4 ensures that factors of irreducible holonomy $H=SU(n_{j})$ or $Sp(\frac{n_{k}^{\prime}}{2})$ can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature.
+Statement (1) of Theorem 1.4 parallels the statement of [9, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved.
+Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([10]), De Rham’s splitting theorem ([17]) and Berger’s classification ([3]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([9, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy.
+Specifically, $Y_{j}$ are Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ are holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ are rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
+Comparing with [9, Theorem 1.4] regarding nonnegative Ricci curvature, Ricci-flat factors can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature.
+As another application of Theorem 1.4, we affirm Problem 1.3 for the simply connected case, which constitutes the third main result of our paper.
 Theorem 1.7.
-Let $(M,\omega)$ be a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature.
-Then $M$ admits a holomorphic and isometric splitting $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ where $M_{l}$ are rationally connected manifolds with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is Hermitian symmetric of compact type).
-If the following splitting theorem for nonnegative holomorphic sectional curvature established, Problem 1.3 will be resolved.
-Question 1.8.
+If $(M,\omega)$ is a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature, then $M$ is projective and rationally connected.
+According to Theorem 1.7, Problem 1.3 is equivalent to the following splitting theorem for nonnegative holomorphic sectional curvature, which is an analogue of Cheeger-Gromoll’s splitting theorem ([10]) for nonnegative Ricci curvature.
+Problem 1.8.
 Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature.
 Can the universal cover $\widetilde{M}$ of $M$ splits isometrically as $N\times\mathbb{C}^{k}$ for some compact Kähler manifold $N$?
-For the Hermitian category, Yang ([36, Corollary 1.10]) proved that a compact Kähler surface admits a Hermitian metric with positive holomorphic sectional curvature is projective and rationally connected.
-In [38], Yang-Zheng introduced the concept of real bisectional curvature as in the following way and generalized Wu-Yau’s Theorem ([19, 34, 35, 16]) to the Hermitian case.
-The underlying reason for the extension is that the Schwarz lemma works the same when the target metric is only Hermitian but has nonpositive real bisectional curvature on a Kähler background.
-Also, they proposed the following generalization of Yau’s Conjecture 1.1 for the Hermitian case ([38, Conjecture 1.6 (b)]).
+Matsumura has affirmed Problem 1.3 for compact Kähler surfaces ([30, Corollary 1.5]).
+To support Problem 1.3 and 1.8 in dimension 3, we provide the following classification of non-projective Kähler 3-dimensional manifolds with nonnegative holomorphic sectional curvature by considering Albanese maps.
+This classification can also serve as a concrete example.
+Theorem 1.9.
+Let $(M,\omega)$ be a non-projective compact Kähler 3-dimensional manifold with nonnegative holomorphic sectional curvature.
+Then $M$ is either flat or a $\mathbb{P}^{1}$-bundle over a torus $\mathbb{T}^{2}$.
+In the latter case, the universal cover of $M$ is $\mathbb{P}^{1}\times\mathbb{C}^{2}$ endowed with the product of a Kähler metric with quasi-positive holomorphic sectional curvature on $\mathbb{P}^{1}$ and the flat metric on $\mathbb{C}^{2}$.
+Remark 1.10.
+From Proposition 4.2 that was used in the proof of Theorem 1.9, it is likely that there exists a parallel result similar to [16, Proposition 3.10] for compact Kähler manifolds with nonnegative holomorphic sectional curvature.
+If this is the case, then Problem 1.3 could be resolved by considering Albanese maps as discussed in [16, Section 3].
+However, we currently do not have any information about the fundamental group of such manifolds.
+Therefore, from this perspective, it might be more necessary and direct to consider Problem 1.8.
+Finally, we consider the Hermitian case.
+Yang ([42, Corollary 1.10]) proved that a compact Kähler surface with a Hermitian metric of positive holomorphic sectional curvature is projective and rationally connected.
+It remains difficult to consider the Hermitian case in dimension $\geq$ 3.
+In [44], Yang and Zheng introduced the concept of real bisectional curvature and generalized Wu-Yau’s Theorem ([22, 40, 41, 18]) to the Hermitian case.
+This concept is also referred as ”positive curvature operator” by Lee and Streets ([28]).
+Additionally, Yang and Zheng proposed a generalization of Yau’s Conjecture 1.1 for the Hermitian case ([44, Conjecture 1.6 (b)]).
 Conjecture 1.11.
 Let $(M,h)$ be a Hermitian manifold with positive real bisectional curvature. Then M is a projective and rationally connected manifold.
-We say that a Hermitian manifold $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
+We say that $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
 $B_{h}(e,a)=\frac{1}{|a|^{2}}\sum\limits_{i,j=1}^{n}R_{i\bar{i}j\bar{j}}a_{i}a_{j}>0,$
 where $R$ is the Chern curvature tensor of $h$.
-The term $B_{h}\geq 0,\ B_{h}<0$ or $B_{h}\leq 0$ are defined similarly.
-This curvature condition is the same as positive holomorphic sectional curvature when the metric $h$ is Kähler and slightly stronger in the non-Kähler case (cf.
+This curvature condition is equivalent to positive holomorphic sectional curvature when the metric $h$ is Kähler, and it’s slightly stronger in the non-Kähler case (cf.
 [44]).
-Under the entra assumption that $M$ is Kählerian, Conjcture 1.9 is just a weaker version of Conjecture 1.2.
-In this special case, Tang ([33, Theorem 1.4]) confirmed the Conjecture 1.9 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$.
-Since the curvature term of Bochner-type formula in the proof of Proposition 3.1 can be controlled by real bisectional curvature, we confirm Conjecture 1.9 under the additional assumption that $M$ is Kählerian.
+Under the entra assumption that $M$ is Kählerian, Conjcture 1.11 is just a weaker version of Conjecture 1.2.
+In this special case, Tang ([39, Theorem 1.4]) confirmed the Conjecture 1.11 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$.
+Applying the Bochner-type formula used in the proof of Proposition 3.1, we confirm Conjecture 1.11 under the additional assumption that $M$ is Kählerian.
 Theorem 1.12.
@@
 In Section 2, we recall some basic definitions and facts.
-In section 3, we give a proof of Theorem 1.4 by establishing an integral inequality.
-In section 4 , we give a proof of Theorem 1.5 and Theorem 1.7.
+In Section 3, we establish an integral inequality to show Theorem 1.4 and provides the proofs of Theorem 1.5 and Theorem 1.7.
+In section 4, we present a proof of Theorem 1.9.
 In Section 5, we give a proof of Theorem 1.12.
+In the appendix, we provide a proof of $L^{2}$-vanishing theorem.
 2. Preliminaries
-For reader’s convenience, we gather some basic definitions and useful facts known to experts, which will be referenced later in the text.
+For the reader’s convenience, we gather some basic definitions and useful facts relevant to the study of rational connectedness and holomorphic sectional curvature, which are well-known to experts and will be referenced later in the text
 2.1. Pseudo-effective line bundle and plurisubharmonic function
@@
 We say that $L$ is pseudo-effective if $L$ can be equipped with a singular Hermitian metric $h$ with $T=\frac{\sqrt{-1}}{2\pi}\Theta_{L,h}=\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ as a current, equivalently, the weight $\varphi\in Psh(\Omega)\cap L_{loc}^{1}(\Omega)$ in any trivalization.
-Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [13, Chapter 1 and Chapter 3]).
+Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [14, Chapter 1 and Chapter 3]).
 Lemma 2.3.
@@
 Lemma 2.4.
-([3, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
+([4, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
 $\lim\limits_{j\rightarrow\infty}v_{j}^{i}=v^{i}\in Psh(\Omega)\cap L^{\infty}_{loc}(\Omega),\ 1\leq i\leq k.$
@@
 2.2. Some criterions for rational connectedness
-Based on the uniruledness criterion estabished in ([6]), Campana-Demailly-Peternell establish the following criterions for rational connectedness.
-Lemma 2.5 ([8, Theorem 1.1]).
+Based on the uniruledness criterion estabished in ([7]), Campana, Demailly and Peternell devoloped the following criterion for rational connectedness.
+Lemma 2.5 ([9, Theorem 1.1]).
 Let $M$ be a projective manifold. The following properties are equivalent.
@@
   3. (c)
-    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(T^{1,0}M^{*})^{\otimes p})$, $p\geq 1$, $\mathcal{F}$ is not pseudo-effective.
-  4. (d)
     For any ample line bundle $A$ on $X$, there exists a constant $C_{A}>0$ such that
@@
 Lemma 2.6.
-(Royden’s observation, [32, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors.
+(Royden’s observation, [38, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors.
 If $S(\xi,\bar{\eta},\zeta,\bar{\omega})$ is a symmetric bihermitian form on some complex vector space $V$ (i.e., $S(\xi,\bar{\eta},\zeta,\omega)=S(\zeta,\bar{\eta},\xi,\bar{\omega})$ and $S(\xi,\bar{\omega},\zeta,\bar{\eta})$), such that for all $\xi\in V$, we have
@@
 Lemma 2.7.
-([37, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$.
+([43, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$.
 Let $e_{1}\in T_{x_{0}}^{1,0}M$ be a unit vector which minimizes the holomorphic sectional curvature $H$ of $\omega$ at $x_{0}$, then
@@
 Lemma 2.8.
-([20, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$.
+([23, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$.
 Assume that the holomorphic sectional curvature $H$ of $\omega$ is nonnegative at $x_{0}$.
@@
 for any $X,Y,Z\in T_{x_{0}}^{1,0}M$, i.e. $V$ is a nonzero truly flat tangent vector at $x_{0}$.
-3. Proof of Theorem 1.4
-In this subsection, we prove Theorem 1.4.
-The proof is divided into two parts: establishing an integral inequality and performing a pointwise computation.
+2.4. The structure theorem of morphism
+The following theorem reveals a detailed structure of morphisms whose domain has semi-positive holomorphic sectional curvature.
+Lemma 2.9.
+[31, Theorem 1.6] Let $(X,g)$ be a compact Kähler manifold with semi-positive holomorphic sectional curvature, and let $\varphi:X\to Y$ be a surjective morphism to a compact Kähler manifold $Y$ with a pseudo-effective canonical bundle.
+Then
+  1. (1)
+    The following statements hold:
+    1. The morphism $\varphi:X\to Y$ is smooth and locally trivial.
+    2.
+The exact sequence of vector bundles $0\to T_{X/Y}\to TX\xrightarrow{d\varphi^{*}}\varphi^{*}TY\to 0$ admits a holomorphic orthogonal splitting, that is, the orthogonal complement $T_{X/Y}^{\perp}$ is a holomorphic vector bundle and there exists an isomorphism $j:\varphi^{*}TY\to T_{X/Y}^{\perp}$ such that it gives the holomorphic orthogonal decomposition
+      $TX=T_{X/Y}\oplus j(\varphi^{*}TY)\cong T_{X/Y}\oplus\varphi^{*}TY.$
+    3.
+The image $Y$ admits a Kähler metric $g_{Y}$ such that $g_{Q}=\varphi^{*}g_{Y}$ and the holomorphic sectional curvature of $g_{Y}$ is identically zero.
+Here $g_{Q}$ is the Hermitian metric on $\varphi^{*}TY$ induced by the above exact sequence and the metric $g$.
+  2. (2)
+    Let $F$ be a fiber of $\varphi$ and $F_{\text{univ}}$ be its universal cover. We obtain the isomorphism
+    $X_{\text{univ}}\cong\mathbb{C}^{m}\times F_{\text{univ}},$
+    where $m$ is the dimension of $Y$.
+There exists a Kähler metric $g_{F_{\text{univ}}}$ on $F_{\text{univ}}$ such that the holomorphic sectional curvature of $g_{F_{\text{univ}}}$ is semi-positive and the above isomorphism $X_{\text{univ}}\cong\mathbb{C}^{m}\times F_{\text{univ}}$ is isometric with respect to the Kähler metrics induced by $g$, $g_{Y}$, and $g_{F_{\text{univ}}}$.
+3. Proof of Theorem 1.4, Theorem 1.5 and Theorem 1.7
+In this section, we provide the proofs of Theorem 1.4, Theorem 1.5 and Theorem 1.7.
+First, we present a proof of Theorem 1.4.
+The proof of Theorem 1.4 is divided into two parts: establishing an integral inequality and performing a pointwise computation.
 The method is based on integration by parts, Bochner-type formula and algebraic properties of holomorphic sectional curvature.
 Setting.
-Let $(M,\omega)$ be a compact Kähler manifold.
+Let $(M,\omega_{g})$ be a compact Kähler manifold.
 Assume that $\mathcal{F}$ is an invertible sheaf of $\mathcal{O}(\Lambda^{p,0}M)$ which is pseudo-effective as a line bundle.
-Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the Čech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$.
-We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([8]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|_{g}}{|\eta_{\alpha}|_{h}}$.
-Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current,
-$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha})$
-is also a closed positive current on $U_{\alpha}$.
+Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the C̆ech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$.
+We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([9]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|^{2}_{g}}{|\eta_{\alpha}|^{2}_{h}}$.
+Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current, We have
+$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}),$
+which is also a closed positive current on $U_{\alpha}$.
 This shows that $\partial\varphi_{\alpha}$ is $L^{2}$ with respect to the wight $e^{\varphi_{\alpha}}$ and $e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}$ has locally finite measure corfficients on $U_{\alpha}$.
@@
 3.1. Proof of the integral inequality (3.1)
-This subsection is devoted to the proof of Proposition 3.1. Firstly, we will proof the following key identity for forms.
+This subsection is devoted to the proof of Proposition 3.1. Firstly, we will show the following key identity for forms.
 Lemma 3.2.
@@
 at $x_{0}$. So, we have
-$\omega^{n-p-1}$
-$=$ $(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}},$
+$\omega^{n-p-1}=(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}}$
 and
-$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}$
-$=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
+$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
 +(terms involving ηJpηKp¯ with Jpπ(Kp) for some permutation $+(\text{terms involving $\eta_{J_{p}}\overline{\eta_{K_{p}}}$ with $J_{p}\neq\pi(K_{p})$ for some permutation $\pi$}).$
@@
 $\begin{split}n&\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}=e^{\varphi_{\alpha}}\operatorname{tr}_{\omega}(\sqrt{-1}\partial\bar{\partial}\psi)|\eta_{\alpha}|^{2}_{g}\cdot\omega^{n}\\ =&p\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}\end{split}$ (3.9)
-where $\beta_{\alpha}=\Lambda^{p-1}(\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{(p-1)!})$.
+where $\beta_{\alpha}=\Lambda^{p-1}((\sqrt{-1})^{p^{2}}\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{p!})$.
 $e^{\varphi_{\alpha}}\beta_{\alpha}$ and $e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}$ are globally defined differential forms with bounded measurable cofficients.
@@
 Step 1. The integration by parts
-Although $\varphi_{\alpha}$ is in general singular, we can still do the integration by parts as in [12].
-For reader’s convenience, we give a detailed explaination.
-Take a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$.
+Although $\varphi_{\alpha}$ is in general singular, we can still perform integration by parts as in [13]. For the reader’s convenience, we provide a detailed explaination.
+Consider a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$.
 Since $\mathrm{supp}\rho_{\alpha}\subset\subset U_{\alpha}$, for any $\epsilon>0$ sufficiently small, there exists a sequence of smooth plurisubharmonic functions $\varphi_{\alpha,\epsilon}$ decreasing to $\varphi_{\alpha}$ on some open subset $V_{\alpha}$ such that $\mathrm{supp}\rho_{\alpha}\subset\subset V_{\alpha}\subset\subset U_{\alpha}$.
-Now the results of Bedford-Taylor (Lemma 2.4) applied to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, imply that we have local weak covergence
+Applying the results of Bedford-Taylor (Lemma 2.4) to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, we obtain local weak covergence
 $e^{\varphi_{\alpha,\epsilon}}\partial\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha},\ e^{\varphi_{\alpha,\epsilon}}\rightarrow e^{\varphi_{\alpha}}\ \text{and}\ e^{\varphi_{\alpha,\epsilon}}\partial\varphi_{\alpha,\epsilon}\wedge\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}$
-on $V_{\alpha}$, then
+on $V_{\alpha}$. Thus
 $\begin{split}&-\int_{M}\left|\partial\psi\right|^{2}_{g}\cdot\omega^{n}=-n\int_{M}\partial\psi\wedge\bar{\partial}\psi\wedge\sum\limits_{\alpha}\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\sum\limits_{\alpha}\int_{V_{\alpha}}e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\int_{M}\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}.\end{split}$ (3.10)
-The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combine with (3.9), we get
+The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combining (3.10) with (3.9), we obtain
 $\begin{split}0\geq&-\int_{M}|\partial\psi|_{g}^{2}\cdot\omega^{n}\\ =&p\int_{M}\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\int_{M}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}.\end{split}$ (3.11)
-Next, let us separately analysis each term on the right-hand side of integral inequality (3.11). Firstly, we consider the sign of first term of (3.11).
+Next, we analyze each term on the right-hand side of integral inequality (3.11) seperately. First, we consider the sign of first term in (3.11).
 Step 2. The Bochner formula.
-For any $x_{0}\in U_{\alpha}$, we choose local holomorphic coordinates centered at $x_{0}$ such that
+For any $x_{0}\in U_{\alpha}$, choose local holomorphic coordinates centered at $x_{0}$ such that
 $\omega=\sqrt{-1}\sum\limits_{i}dz^{i}\wedge d\bar{z}^{i}$
@@
 $\beta_{\alpha}=\sqrt{-1}\sum\limits_{i}\beta_{i\bar{i}}dz^{i}\wedge d\bar{z}^{i}$
-at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$, recall (3.3) and (3.7) that we have
+at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$. Recall (3.3) and (3.7), which give us
 $\operatorname{tr}_{\omega}\beta=|\eta|_{g}^{2}$ (3.12)
@@
 $\beta_{i\bar{i}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\geq 0.$ (3.13)
-By a direct calculation, we derive the following Bochner type formula in the sense of distribution,
+By direct calculation, we derive the following Bochner type formula in the sense of distribution,
 $\begin{split}\partial_{u}\partial_{\bar{u}}(e^{\varphi}|\eta|^{2}_{g})=&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p!\sum\limits_{I_{p}}\sum\limits_{k=1}^{p}\sum\limits_{l=1}^{n}R_{u\bar{u}i_{k}\bar{l}}\eta_{I_{p}}\overline{\eta_{i_{1}\cdots(l)_{k}\cdots i_{p}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p(p!)\sum\limits_{l,k=1}^{n}R_{u\bar{u}l\bar{k}}\sum\limits_{J_{p-1}}\eta_{lJ_{p-1}}\overline{\eta_{kJ_{p-1}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p\sum\limits_{l,k}R_{u\bar{u}l\bar{k}}\beta_{l\bar{k}})\end{split}$ (3.14)
@@
 It follows from Royden’s observation(Lemma 2.6) that
-$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=&e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\\ \geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
+$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
 at $x_{0}$. The fact that $\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ and $\beta\geq 0$ implies that
@@
 As
-¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
+¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+e^{\varphi_{\alpha}}(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
 $=$ $(\sqrt{-1})^{2p}e^{\varphi_{\alpha}}(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})\wedge\overline{(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})}+e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}},$
@@
 Since $\varphi_{\alpha}\in Psh(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$, $\{\varphi_{\alpha}=-\infty\}$ has zero Lebesgue measure.
-This fact implies that $\eta_{\alpha}=0$ almost everywhere on $U_{\alpha}$.
-Then $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$.
-Therefore, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$.
+This fact implies that $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts the fact that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$.
+Hence, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$.
 The following calculation will be done in the smooth sense.
@@
 Since $x_{0}\in M$ is arbitrary, we have ¯$\partial\bar{\partial}\varphi=0$ on $M$ and $D^{\prime}\eta+\eta\partial\varphi=0$ on $M$.
-¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle.
-In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, then $\varphi_{\alpha}$ can be written $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$ and
-$D^{\prime}(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$
+The fact that ¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle.
+In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, so it can be written as $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$.
+Therefore,
+$D(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$ (3.25)
 where $e^{f_{\alpha}}\eta_{\alpha}$ is a local holomorphic section of $\mathcal{F}|_{U_{\alpha}}$.
-This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$.
-The proof is completed.
+This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$, which verifies the statement (1).
+We obtain a global real semi-positive $(1,1)$-form given by
+$\beta:=e^{\varphi_{\alpha}}\beta_{\alpha}=\Lambda^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{e^{f_{\alpha}\eta_{\alpha}}\wedge\overline{e^{f_{\alpha}}\eta_{\alpha}}}{p!}\right),$
+which is determined by $\mathcal{F}$ and is parallel with respect to $D$. We then define a parallel and self-adjoint linear transformation $S:T^{1,0}M\rightarrow T^{1,0}M$ by
+$\langle S(X),Y\rangle_{g}=\beta(X,\bar{Y}),\ \forall X,Y\in T^{1,0}M.$ (3.26)
+From the above computation at a fixed point $x_{0}$, $V_{x_{0}}=\operatorname{span}\{\frac{\partial}{\partial z^{i}}\}_{i=1}^{m}$ and $W_{x_{0}}=\operatorname{span}\{\frac{\partial}{\partial z^{i}}\}$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S_{x_{0}}$, respectively.
+We also have that every $v\in T_{x_{0}}^{1,0}M$ is truly flat and $\dim V_{x_{0}}=m>0$.
+Since $S$ is parallel, $x\mapsto V_{x}$ and $x\mapsto W_{x}$ define parallel distrubutions.
+The proof is complete.
 ∎
-4. Proof of Theorem 1.5 and Theorem 1.7
-In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7.
-The proof of Theorem 1.5 is based on Theorem 1.4 (1) and Campana-Demailly-Peternell’s critetion.
-The proof of Theorem 1.7 follows the arguement in [7, section 4] for $\operatorname{Ric}\geq 0$, demonstrating that factors of irreducible holonomy $H=U(n)$ are rationally connected manifolds.
-The difference is that factors of irreducible holonomy $H=SU(n)$ or $Sp(\frac{n}{2})$ can be excluded by statement (1) of Theorem 1.4.
+3.3. Proof of Theorem 1.5 and Theorem 1.7
+In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7. The proofs are based on Theorem 1.4, Campana-Demailly-Peternell’s critetion, de Rham’s splitting theorem ([17]).
 Proof of Theorem 1.5.
-Since the holomorphic sectional curvature is nonnegative and there exists no nonzero truly flat tangent vector, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective by the statement (1) of Theorem 1.4.
-This fact implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$, in particular, M is projective by Kodaira’s theorem ([21]).
-Combine with Campana-Demailly-Peternell’s criterion (Lemma 2.5), $M$ is rationally connected.
+Since the holomorphic sectional curvature is nonnegative and there is no nonzero truly flat tangent vector at some point, by Theorem 1.4 any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective.
+This actually implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$.
+In particular, M is projective by Kodaira’s theorem ([25]).
+Combining this with Campana-Demailly-Peternell’s criterion (Lemma 2.5), we conclude that $M$ is rationally connected.
 ∎
 Proof of Theorem 1.7.
-Let $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ be the De Rham decomposition of $(M,\omega)$, induced by a decomposition of the holonomy representation in irreducible representations ([4]).
-Since the holonomy is contained in $U(n)$, all factors $(M_{l},\omega_{l})$ are Kähler manifolds with irreducible holonomy and holonomy group $H_{l}\subset U(n_{l})$, $n_{l}=\dim M_{l}$.
-The product structure implies that $\omega_{l}$ has nonnegative holomorphic sectional curvature.
-By Berger’s classification of holonomy groups ([4]), there are only three possibilities: $H_{l}=U(n_{l}),\ SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$, unless $M_{l}$ is a Hermitian symmetric space, then necessarily of compact type (such symmetric space were classified by Élie Cartan, they are rational homogeneous, hence rationally connected, their holonomy groups are also known, see.
-e.g.
-[5]).
-Case 1: $H_{l}=SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$.
-Then $M_{l}$ is a Calabi-Yau manifold or a holomorphic symplectic manifold (see e.g.
-[1]).
-Recall that a simply connected Calabi-Yau manifold $M_{l}$ is a compact simply connected projective manifold with $K_{M_{l}}\simeq O_{M_{l}}$ and $H^{0}(M_{l},\Omega_{M_{l}}^{p})=0$ for $0<p<\dim M_{l}$.
-A holomorphic symplectic manifold is a compact Kähler manifold $X$ which admits a holomorphic symplectic $(2,0)$-form (of maximal rank everywhere).
-In particular, $M_{l}$ admits a holomorphic $(n_{l},0)$-form that vanishes nowhere.
-Recall that we can define a nonnegative $(1,1)$-form $\beta=\Lambda^{p-1}({\frac{\eta\wedge\overline{\eta}}{n_{l}!}})$.
-For any point $x_{0}$ of $M_{l}$, we can choose some local holomorphic coordinates such that the coefficients of $w_{l}$ is identity and $\beta$ is diagonal at $x_{0}$.
-Locally, we can write $\eta=\frac{1}{n_{l}!}\sum\limits fdz^{1}\wedge\cdots dz^{n}$.
-It follows from the local representation (3.7) of $\beta$ that we have
-$\beta=\frac{|\eta|^{2}}{n_{l}}\sum\limits_{i=1}^{n_{l}}dz^{i}\wedge d\overline{z}^{i}$
-at $x_{0}$.
-The proof of Theorem 1.4 tells us that $\{\frac{\partial}{\partial z_{i}}\}_{i=1}^{n_{l}}$ are truly flat tangent vectors at $x_{0}$.
-Hence the Riemannian curvature tensor $\operatorname{Rm}$ of $\omega_{l}$ vanishes at $x_{0}$.
-Since $x_{0}$ is arbitrary, $\operatorname{Rm}$ vanishes everywhere, $M_{l}$ is a simply connected complete manifold with zero holomorphic sectional curvature.
-By classification of complex space forms, $M_{l}$ must be holomorphically isometric to complex Euclidean space $\mathbb{C}^{n_{l}}$, which contradicts the fact that $M_{l}$ is compact.
-Case 2: $H_{l}=U(n_{l})$.
-Since the natural representation of $U(n_{l})$ on $(\mathbb{C}^{n_{l}})^{\otimes p}$ has no invariant line on which $U(n_{l})$ induces a trivial action, there cannot exist any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M_{i})$, $1\leq p\leq n$ such that $\mathcal{F}$ is invariant by parallel transport induced by the connection $D$.
-Combine with statement (2) of Theorem 1.4, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective.
-Therefore $M_{l}$ is rationally connected by Campana-Demailly-Peternell’s criterion.
-In summary, each $(M_{l},\omega_{l})$ is a rationally connected manifold with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is a hermitian symmetric space of compact type).
+We will argue by contradiction.
+Suppose there exists an invertible subsheaf $\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}M)$ that is pseudoeffective as a line bundle.
+According to Theorem 1.4, $T^{1,0}M$ admits a decomposition $V+W$ of $T^{1,0}M$ where $V$ and $W$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S$, respectively, and $\forall p\in M$, every $v\in V_{p}$ is truly flat.
+Since $M$ is simply connected, by de Rham’s splitting theorem ([17, Theorem II]), $(M,\omega)\simeq(M_{1},\omega_{1})\times(M_{2},\omega_{2})$ isometrically and holomorphically, with $TM_{1}=V$.
+This implies that $M_{1}$ is a simply connected Kähler manifold with the flat metric $\omega_{1}$.
+According to the classification of complex space forms, $M_{1}$ must be $\mathbb{C}^{m}$ with $m=\dim V>0$, which contradicts the fact that $M$ is compact.
+By Campana-Demailly-Peternell’s criterion (Lemma 2.5), the proof is complete.
 ∎
 4. Proof of Theorem 1.9
-In this section, we give a proof of Theorem 1.10.
+In this section, we provide a proof of Theorem 1.9.
+The proof relies on $L^{2}$-vanishing theorem (Proposition 4.1), Atiyah’s $L^{2}$-index Theorem ([1]) and properties of Albanese maps.
+First, we present the following $L^{2}$-vanishing theorem for nonnegative holomorphic sectional curvature.
+Proposition 4.1.
+Let $(M,\omega)$ be a complete Kähler manifold with nonnegative holomorphic sectional curvature, then any $L^{2}$-holomorphic $(p,0)$-form is parallel.
+The idea of proof is essentially the same as that for Theorem 1.4, but with some minor differences, because $|\eta|^{4}_{g}$ may be not $L^{1}$-integrable.
+For the reader’s convenience, we provide a brief proof in the appendix.
+From Proposition 4.1 and Atiyah’s $L^{2}$-index theorem ([1]), we derive the following proposition.
+Proposition 4.2.
+Let $(M,\omega)$ be a compact Kähler manifolds with nonnegative holomorphic sectional curvature. One of the two following situations occurs:
+  • (i)
+    $\pi_{1}(M)$ is finite, then $M$ is projective and rationally connected.
+  • (ii)
+    $\pi_{1}(M)$ is infinite, then $\chi(M,\mathcal{O}_{X})=0$.
+Proof.
+Let $f:\widetilde{M}\rightarrow M$ be the universal cover of $M$, then $(\widetilde{M},f^{*}\omega)$ is a compact simple connected Kähler manifold with nonnegative holomorphic sectional curvature.
+If $\pi_{1}(M)$ is finite, then $(\widetilde{M},f^{*}\omega)$ is compact simple connected, hence projective and rationally connected (Theorem 1.7).
+Since a pseudoeffective invertible subsheaf $\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}M)$ induces a pseudoeffective invertible subsheaf $f^{*}\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}\widetilde{M})$, $M$ must be projective and rationally connected by the Compana-Demailly-Peternell’s criterion (Lemma 2.5).
+If $\pi_{1}(M)$ is infinite, then $(\widetilde{M},f^{*}\omega)$ is noncompact and complete.
+Denote the Hilbert space of $L^{2}$-harmonic form of degree $(p,q)$ by $H^{p,q}_{L^{2}}(\widetilde{M})$.
+According to $L^{2}$-Hodge theory (see [20]), a $L^{2}$ $(p,0)$-form is harmonic if and only if it is a holomorphic $(p,0)$-form.
+In general, $H^{p,q}_{L^{2}}(\widetilde{M})$ might be infinite-dimensional.
+Using the isometric action of $\Gamma:=\pi_{1}(X)$ on them, one can associate to them a nonnegative real number $\dim_{\Gamma}(H^{p,q}_{L^{2}}(\widetilde{M}))$ (c.f.
+[1]).
+$\dim_{\Gamma}(H^{p,q}_{L^{2}}(\widetilde{M}))=0$ if and only if $H^{p,q}_{L^{2}}(\widetilde{M})=\{0\}$.
+By Atiyah’s $L^{2}$-index theorem (c.f.
+[1, 20]), we know that
+$\chi(M,\mathcal{O}_{M})=\chi_{L^{2}}(\widetilde{M},\mathcal{O}_{\widetilde{M}}):=\sum\limits_{k=0}^{\dim M}(-1)^{k}\dim_{\Gamma}(H^{0,q}_{L^{2}}(\widetilde{M})).$
+By Proposition 4.1, the norm $|\eta|_{f^{*}g}$ of any $L^{2}$-holomorphic $(p,0)$-form $\eta$ is equal to some constant.
+Because the volume of $(\widetilde{M},f^{*}\omega)$ is infinite, any $L^{2}$-holomorphic $(p,0)$-form must be zero.
+Thus $H^{p,0}_{L^{2}}(M)=\{0\}$ for every $0\leq p\leq n$.
+By conjugation, $H^{0,p}_{L^{2}}(M)=\{0\}$ for every $0\leq p\leq n$.
+Consequently, $\chi(M,\mathcal{O}_{M})=0$.
+∎
+From Proposition 4.1, we have the following fact about the Albanese map.
+Proposition 4.3.
+Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Then the Albanese map $\alpha:=M\rightarrow A(M)$ is a submersion. In particular, it is surjective.
+Proof.
+Let $q=\dim H^{1,0}_{\bar{\partial}}(M)$ and let $\{\eta_{1},\cdots,\eta_{q}\}$ be a basis of holomorphic $(1,0)$-forms. Recall that the Albanese map $\alpha$ is defined by
+$x\mapsto\left(\int_{a}^{x}\eta_{1},\cdots,\int_{a}^{x}\eta_{q}\right)\mod\Lambda,$
+where $\Lambda\subset\mathbb{C}^{q}$ (see, e.g., [14, Chapter VI, section 9.2]).
+Then $d\alpha=(\eta_{1},\cdots,\eta_{q})$ at every point.
+If $\operatorname{rank}(d\alpha)<q=\dim(A(M))$ at some point $x$, then there would exist a non-zero linear combination
+$u=\lambda_{1}\eta_{1}+\cdots+\lambda_{q}\eta_{q}$
+with $u(x)=0$. By Proposition 4.1, $u$ must be zero, which leads to a contradiction. ∎
+Now let us proceed to show Theorem 1.9.
+Proof of Theorem 1.9.
+Since $M$ is non-projective, it follows from Kodaira’s theorem ([25]) that $M$ admits a nonzero holomorphic $(2,0)$-form, and from Proposition 4.2 that $\chi(M,\mathcal{O}_{M})=0$.
+Let $h^{p,0}$ denote $\dim H^{p,0}_{\bar{\partial}}(M)$.
+If $h^{3,0}>0$, there admits a nonzero holomorphic $(3,0)$-form $\eta$.
+By Proposition 4.1, $\eta$ must be parallel and therefore doesn’t vanish anywhere.
+Since $\dim M=3$, the existence of the nowhere-vanishing holomorphic $(3,0)$-form $\eta$ implies that $K_{M}\simeq\mathcal{O}_{M}$.
+Consequently, $K_{M}$ is flat, and we have:
+$0=-n\int_{M}c_{1}(K_{M})\wedge\omega^{n-1}=n\int_{M}\operatorname{Ric}(\omega)\wedge\omega^{n-1}=\int_{M}S\cdot\omega^{n},$
+where $S$ is the scalar curvature of $\omega$. According to Berger’s averaging trick ([3]),
+[…]
+at any point $x$ of $M$. Hence, the holomorphic sectional curvature is identically zero, and thus $M$ is flat.
+If $h^{3,0}=0$, then $01-h^{1,0}+h^{2,0}=\chi(M,\mathcal{O}_{M})=$ 0, which implies that $h^{1,0}=1+h^{2,0}\geq 2$.
+Suppose that $h^{1,0}=3$.
+Then we have $h^{3,0}\geq 1$, which is a contradiction.
+Therefore, $h^{1,0}=2$.
+From Matsumura’s result (Lemma 2.9) and Proposition 4.3, the Albanese map $\alpha:M\rightarrow\mathbb{T}^{2}$ is locally trivial and smooth.
+The fibre $F$ of $\alpha$ admits a Kähler metric $\omega_{F}$ with nonnegative holomorphic sectional curvature, and the universal cover $(\widetilde{M},\widetilde{\omega})\cong(\widetilde{F},\widetilde{\omega_{F}})\times(\mathbb{C}^{2},\omega_{0}),$ where $(\widetilde{M},\widetilde{\omega})$ and $(\widetilde{F},\widetilde{\omega_{F}})$ are the universal covers of $(M,\omega)$ and $(F,\omega_{F})$ respectively, $\omega_{0}$ is the flat metric.
+Since $\dim F=1$, the holomorphic sectional curvature of $\omega_{F}$ is either identically zero or quasi-positive.
+Hence, either $M$ is flat, or $F$ is $\mathbb{P}^{1}$.
+When $F$ is $\mathbb{P}^{1}$, $M$ is a $\mathbb{P}^{1}$-bundle over a torus $\mathbb{T}^{2}$, and the universal cover is $\mathbb{P}^{1}\times\mathbb{C}^{2}$.
+∎
+5. Proof of Theorem 1.12
+In this section, we provide a proof of Theorem 1.12.
 Proof of Theorem 1.12.
@@
 By a direct computation, we have the following Bochner type formula,
-$\partial_{u}\partial_{\bar{u}}|\psi|^{2}$
-$=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
+$\partial_{u}\partial_{\bar{u}}|\psi|^{2}=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
 $-\sum\limits_{t,s=1}^{r}R^{H}_{u\bar{u}t\bar{s}}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,s)}}$
@@
 with $\iota^{p+k}_{(e^{t})^{*}}(\psi)=\psi(\cdots;\cdots,(e^{t})^{*},\cdots)$ where the contractions are done at the $(p+k)$-th slot.
-Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then
-$\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=\sum\limits_{i=1}^{n}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,i),J_{l}}}=p|\psi|^{2}$ (5.5)
-and
-$\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=\sum\limits_{t=1}^{r}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,t)}}=l|\psi|^{2}.$ (5.6)
-So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,
+Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then $\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=p|\psi|^{2}$ and $\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=l|\psi|^{2}.$ So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,
 $\langle\sqrt{-1}\partial\bar{\partial}|\psi|^{2},\alpha\rangle_{h}=$ $\sum\limits_{i=1}^{n}\partial_{i}\partial_{\bar{i}}|\psi|^{2}\alpha_{i\bar{i}}$
@@
 ∎
+Appendix. Proof of $L^{2}$-vanishing Theorem
+Proof of Proposition 4.1.
+Since $\omega$ is complete, there exists an exhaustive sequence $(K_{v})_{v\in\mathbb{N}}$ of compact subsets of $M$ and functions $\psi_{v}\in C^{\infty}(M,\mathbb{R})$ (see.
+e.g.
+[14, Section 8.2]) such that
+  • $\psi_{v}=1$ in a neighborhood of $K_{v}$, $\mathrm{supp}\psi_{v}\subset K^{\circ}_{v+1}$,
+  • $0\leq\psi_{v}\leq 1$ and $|d\psi_{v}|^{2}_{g}\leq 2^{-v}\psi_{v}$ on $M$.
+Let $\eta$ be a nonzero $L^{2}$-holomorphic $(p,0)$-form, then $|\eta|\in W^{1,\infty}_{loc}(M)$ and $\Sigma:=\{x\in M:\eta(x)=0\}$ is a proper subvariety of $M$, from Lemma 3.2 and Bochner-type formula (3.14), we have
+$\begin{split}&-\int_{M}\langle\partial\psi_{v},\bar{\partial}|\eta|_{g}^{2}\rangle_{g}\cdot\omega^{n}=\int_{M}\psi_{v}\Delta|\eta|_{g}^{2}\cdot\omega^{n}\\ \geq&\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\int_{M}\psi_{v}\cdot\kappa|\eta|_{g}^{2}\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\\ \geq&\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\end{split}$ (5.6)
+where $\beta=\Lambda^{p-1}(\frac{\eta\wedge\overline{\eta}}{p!})$.
+Recall (3.3) that $\operatorname{tr}_{\omega}\beta=|\eta_{g}|^{2}$, then $|\frac{\beta}{|\eta|_{g}^{2}}|_{g}\leq 1$ and $|\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}|_{g}\leq 1$ on $M\setminus\Sigma$, thus each integral on the RHS of (5.6) makes senses.
+For any fixed point $x_{0}\in M$, we choose local holomorphic coordinates $(z^{1},\cdots,z^{n})$ centered at $x_{0}$ such that $(g_{i\bar{j}})$ is the identity and $(\beta_{i\bar{j}})$ is diagonal, then
+$4\langle\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g},\beta\rangle_{g}=\frac{\partial_{i}|\eta|_{g}^{2}\cdot{\partial}_{\bar{i}}|\eta|_{g}^{2}}{|\eta|_{g}^{2}}\beta_{i\bar{i}}=\frac{\langle D^{\prime}_{i}\eta,\bar{\eta}\rangle\cdot\langle\eta,\overline{D^{\prime}_{i}\eta}\rangle}{|\eta|_{g}^{2}}\beta_{i\bar{i}}\leq\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\beta\rangle.$ (5.7)
+As the $L^{2}$-holomorphic $(p,0)$-form $\eta$ is closed (see [20]), applying Lebesgue’s dominated covergence theorem, we have
+$\begin{split}0=&\int_{M}\partial(\psi_{v}\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1})\\ =&\lim_{\epsilon\rightarrow 0}\int_{M}\partial(\psi_{v}\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}\wedge\omega^{n-p-1})\\ =&\lim_{\epsilon\rightarrow 0}\int_{M}\left(\partial\psi_{v}\wedge\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}+\psi_{v}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}-4\psi_{v}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{|\eta|_{g}^{2}}{(|\eta|_{g}^{2}+\epsilon)^{2}}\eta\wedge\bar{\eta}\right)\wedge\omega^{n-p-1}\\ =&\int_{M}\left(\partial\psi_{v}\wedge\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}+\psi_{v}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}-4\psi_{v}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\right)\wedge\omega^{n-p-1}.\end{split}$ (5.8)
+Applying Lemma 3.2, we have
+$\int_{M}\psi_{v}|\partial|\eta|_{g}|_{g}^{2}\cdot\omega^{n}=\int_{M}\psi_{v}\operatorname{tr}_{\omega}(\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g})\cdot\omega^{n}$
+$=$ $\int_{M}\psi_{v}\langle\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g},\frac{\beta}{|\eta|_{g}^{2}}\rangle_{g}\cdot\omega^{n}+\frac{(\sqrt{-1})^{p^{2}}n!}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}$
+$\leq$ $\frac{1}{4}\left(\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{(\sqrt{-1})^{p^{2}}n!}{(n-p-1)!}\lim_{\epsilon\rightarrow 0}\int_{M}\psi_{v}\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}\wedge\omega^{n-p-1}\right)$
+$=$ $\frac{1}{4}\left(\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\right)$
+$-\frac{(\sqrt{-1})^{p^{2}}n!}{4(n-p-1)!}\int_{M}\sqrt{-1}\partial\psi_{v}\wedge\bar{\partial}|\eta|^{2}_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}$
+$\leq$ $C_{1}\int_{M}|\partial\psi_{v}|_{g}\cdot\big{|}\partial|\eta|_{g}\big{|}\cdot|\eta|_{g}\cdot\omega^{n}$
+for some constant $C_{1}>0$ independent of $v\in\mathbb{N}$. The fact that $|\partial\psi_{v}|_{g}^{2}\leq 2^{-v}\psi_{v}$ and the Schwarz inequality yields
+$\int_{M}\psi_{v}\big{|}\partial|\eta|_{g}\big{|}_{g}^{2}\cdot\omega^{n}\leq 2^{-v}C_{1}||\eta||_{L^{2}},$
+hence $\int_{M}\psi_{v}\big{|}\partial|\eta|_{g}\big{|}_{g}^{2}\cdot\omega^{n}\rightarrow 0$ for $v\rightarrow\infty$, which implies that $|\eta|_{g}\equiv C$ for some constant $C\geq 0$.
+Applying the pointwise computation in the proof of Theorem 1.4, the fact (3.25) tells us that $D\eta=0$.
+The proof is completed.
+∎